### v2

# Singular basins in multiscale systems: tunneling between stable states

Abstract
Real-world systems often evolve on different timescales and possess multiple coexisting stable states. Whether or not a system returns to a given stable state after being perturbed away from it depends on the shape and extent of its basin of attraction. We show that basins of attraction in multiscale systems can exhibit special geometric properties in the form of singular funnels. Although singular funnels are narrow, they can extend to different regions of the phase space and, unexpectedly, impact the system’s resilience to perturbations. Consequently, singular funnels may prevent common dimensionality reductions in the limit of large timescale separation, such as the quasi-static approximation, adiabatic elimination and time-averaging of the fast variables. We refer to basins of attraction with singular funnels as singular basins. We show that singular basins are universal and occur robustly in a range of multiscale systems: the normal form of a pitchfork bifurcation with a slowly adapting parameter, an adaptive active rotator, and an adaptive network of phase rotators.
Complex real-world systems are characterized by multistability, i.e. when started from different initial conditions, the same system can end up with notably different asymptotic behavior. Examples of multistability can be found in epileptic and neuronal models [27, 19, 42, 38], reservoir computers [17], lasers [26, 56, 57], and climate [1, 45, 49], to name just a few. The analysis and control of multistability-related phenomena has therefore been the subject of many studies [43, 14, 37, 22]. An important question in multistable systems concerns the interplay between different stable states and possible transitions between them. For example, in the context of cell differentiation, such transitions correspond to the reprogramming of cells [24]. One notable approach addressing the above challenges is the concept of basin stability [40, 50], which represents the likelihood of reaching a given state (attractor) when starting from a random initial condition. This likelihood can be estimated by the relative volume of its basin of attraction, i.e., the proportion of initial states in the phase space that converge to this attractor. More importantly, the shape of the basins of attraction determines the uncertainty [11] and resilience [23, 30] of a system to perturbations. Therefore, to understand and control a multistable system, it is crucial to understand the geometry of its basins of attraction.
In addition to having multiple stable states, it is also common for real-world systems to evolve on multiple timescales, which gives rise to a variety of nonlinear phenomena [34, 33, 4, 46, 12]. An important example of a multi-scale system is an adaptive dynamical network, where the adaptation is much slower than the node dynamics [5, 6, 16, 39, 20, 2, 54]. Such systems are challenging because they are also high-dimensional and can exhibit a high degree of multistability. The study of basins of attraction in multiscale systems is an area of research that remains largely unexplored. Therefore, uncovering new properties of the basins that arise from the interplay of different timescales is an intriguing and important task.
Useful and widely used techniques have been developed to analyze mathematical models of multiscale systems with two distinct timescales, also known as slow-fast systems [28, 55, 15, 32]. These techniques exploit the limit of infinite ratio between the fast and slow timescales and can be broadly divided into two groups [25, 21, 44, 34]. The first group includes adiabatic elimination and averaging. It eliminates the fast variables by treating them as instantaneous, or by taking into account their effective contribution, while only considering the time evolution of the slow variables in the reduced slow subsystem. The second group includes the quasistatic approximation. It ‘freezes’ the slow variables by treating them as static at each instant, while only considering the time evolution of the fast variables in the reduced fast subsystem.
This Letter uncovers special properties of basins of attraction in slow–fast systems, leading to counter-intuitive phenomena. We show that such basins can contain singular funnels (SFs) in the form of tunnels that extend to different regions of the phase space and become increasingly narrow as the timescale ratio increases. The problem with SFs is that they are eliminated by the quasistatic approximation and adiabatic elimination. Consequently, adiabatic elimination makes it impossible to reach a given stable state from certain regions of the phase space, whereas this is possible in the full system. On the other hand, the quasistatic approximation could mask the presence of multiple basins of attraction entirely. Therefore, one must be cautious when extrapolating the resilience properties of the full system from those of a reduced one, regardless of whether the latter is obtained via quasistatic approximation, adiabatic elimination, or averaging. Due to the vanishing of SFs in the limit of infinite ratio of the timescales, we refer to basins containing SFs as singular basins.
[…]
$\dot{x}(t)$ $=x(\mu-x^{2}),$ (1)
$\dot{\mu}(t)$ $=\varepsilon(-\mu+ax-b),$ (2)
where the dot denotes $d/dt$. The small parameter $0<\varepsilon\ll 1$ is the ratio between the fast timescale of $x$ and the slow timescale of $\mu$. When $\varepsilon=0$, the slow timescale is infinitely slower, or equivalently, the fast timescale is infinitely faster. The parameters $a,b>0$ determine the adaptation rule. Under the condition $a>2\sqrt{b}$, system (12) has two stable equilibria, $e_{1}$ and $e_{3}$, one saddle equilibrium, $e_{2}$, and a SF, as shown in Fig. 1(a); see the supplemental material for more details.
The fast subsystem is obtained by setting $\varepsilon=0$ in system (12), which gives system (1) with static $\mu$. Since it has a unique stable quasistatic equilibrium $x^{*}(\mu)$ that varies continuously with $\mu$ (the solid part of the gray curve $S$ in Fig. 1(a)), we apply the adiabatic elimination procedure. We substitute $x^{*}(\mu)$ into (2) and obtain the reduced slow subsystem
$\frac{d\mu}{d\tau}=f(\mu),$ (3)
where $f(\mu)=-\mu+a\sqrt{\mu H(\mu)}-b$ and $\tau=\varepsilon t$ the slow time. We then recognise that this is a bistable dynamical system with a double-well potential $U(\mu)=-\int f(\mu)d\mu=\mu^{2}/2+b\mu-2a\left(\mu H(\mu)\right)^{3/2}/3$ and the same equilibria as the full system; see Fig. 1(b). After adiabatic elimination, the basins of attraction of the stable equilibria $e_{1}$ and $e_{3}$ in Fig. 1(b) are clearly separated by a single value of the slow variable, $\mu=\mu_{b}$, given by the position of the unstable equilibrium $e_{2}$. If the slow subsystem is started on the left (resp. right) of $\mu_{b}$, it can only converge to $e_{1}$ (resp. $e_{3}$). However, this clear separation of basins by $\mu=\mu_{b}$ is never present in the full system (12) due to the SF, no matter how large the timescale ratio is (or how small $\varepsilon$ is). In particular, the SF allows convergence to $e_{1}$ from both sides of $\mu_{b}$, which is impossible in the reduced slow subsystem. This includes perturbations from $e_{3}$ into the SF along the fast $x$-direction. From the slow subsystem’s perspective, this appears as though there is quantum-like ‘tunnelling’ or ‘teleporting’ from $e_{3}$ to $e_{1}$ (the grey arrow in Fig. 1(b)). In summary, the SF clearly restricts the applicability of the adiabatic elimination, even when the fast subsystem has a unique stable quasistatic equilibrium that varies continuously with $\mu$.
Figure 1: (a) Singular basin with a singular funnel (SF) in a bistable pitchfork normal form (12) with slowly adapting parameter $\mu$. The red (white) region is the basin of attraction of the stable equilibrium $e_{1}$ ($e_{3}$). The boundary of the two basins is given by the stable manifold of the saddle equilibrium $e_{2}$ (blue curves). The SF near $x=0$ shrinks exponentially with $\mu$. (b) The bistable potential of the corresponding slow subsystem after adiabatic elimination of the fast variable $x$ (equivalently, reduction to the stable critical manifold). Parameters: $a=3$ , $b=2$, and $\varepsilon=0.1$. (c)-(d) The same as (a)-(b) but for system (26) with $a=5$ , $b=10$, and $\varepsilon=0.1$.
One might suspect that the failure of a global adiabatic elimination for system (12) is caused by the degeneracy at $(\mu,x)=(0,0)$, where $x^{*}(\mu)$ is stable but non-hyperbolic. To demonstrate that this is not the case, we consider a modified system
$\dot{x}(t)$ $=x\left(\tanh\mu+2-x\right),$ (4)
$\dot{\mu}(t)$ $=\varepsilon(-\mu+ax-b),$
that does not have any degeneracies. As shown in Fig. 1 (c–d), this system exhibits a singular basin with a SF and the same failure of global adiabatic elimination.
The SFs in the paradigmatic models (12) and (26) have very simple geometry. In general, however, SFs can have intricate geometries and can penetrate the (white) basin of the other stable state in complicated ways. We will now demonstrate this for adaptive phase rotators.
The fast-slow adaptive phase rotator has the form [18]
$\dot{\varphi}(t)$ $=\omega+\mu-\sin\varphi,$ (5)
$\dot{\mu}(t)$ $=\varepsilon(-\mu+\eta(1-\sin(\varphi+\alpha))),$ (6)
where $\varphi\in(0,2\pi]$ is a phase variable, $\omega$ is the base oscillator frequency, and $\mu$ accounts for a slow self-adjustment of the oscillator frequency. The fast subsystem (5) with static $\mu$ has a pair of stable and unstable equilibria for $|\omega+\mu|<1$, and a periodic rotation for $|\omega+\mu|>1$  [^2] This corresponds to the case of unboundedly growing $\varphi(t)$ when $\varphi$ considered on a real line $\mathbb{R}$ or, equivalently, rotating on a circle $\mathbb{S}^{1}$ when $\varphi$ considered modulo $2\pi$ phase. At the point $|\omega+\mu|=1$, the fast subsystem undergoes a "saddle-node bifurcation on a periodic rotation" [18, 52]. Hence, the fast subsystem (5) has a single stable attractor for all fixed values of $\mu$, which is an equilibrium for $|\omega+\mu|\leq 1$ and a periodic rotation for $|\omega+\mu|>1$.
The adaptive phase rotator (56) can also be reduced to the form (3) by eliminating the fast variable $\varphi$. This is achieved by averaging over the fast rotations for $|\omega+\mu|>1$ and by adiabatic elimination for $|\omega+\mu|<1$. The corresponding function $f(\mu)$ of the slow subsystem (3) can be calculated explicitly; see the supplemental material for more details:
$f(\mu)=-\mu+\eta\left(1-\left(\mu+\omega\right)\cos\alpha-\Omega(\mu)\right),$ (7)
where
$\Omega(\mu)=\begin{cases}\sqrt{1-\left(\mu+\omega\right)^{2}}\cdot\sin\alpha,\quad\left|\mu+\omega\right|\leq 1\\ \sqrt{\left(\omega+\mu\right)^{2}-1}\cdot\cos\alpha,\quad\left|\mu+\omega\right|>1.\end{cases}$
We consider the case when system (56) possesses two coexisting attractors: a stable equilibrium $e_{1}$ and a stable rotation $\gamma_{c}$; see Fig. 2(a) and the supplemental material for more details. The slow subsystem (3), with $f(\mu)$ given by equation (7), is also a bistable system with a double-well potential as shown in Fig 2(b).
Figure 2: (a,c) Singular basin with a singular funnel (SF) in an adaptive phase-oscillator (56). The basin of attraction of the periodic rotation $\gamma_{c}$ is plotted in red and the basin of attraction of the stable equilibrium $e_{1}$ is plotted in white. The (blue) boundary of the two basins is given by the stable manifold of the saddle equilibrium $e_{2}$. (b) The bistable potential of the corresponding slow subsystem after adiabatic elimination and averaging of the fast variable $\varphi$. (d) The volume of the SF limited to $10\leq\mu\leq 3$ as a function of $\varepsilon$ obtained using equation (8) (black curve) and Monte Carlo simulation (red open circles) see the supplemental material for details. Other parameters: $\eta=10$, $\omega=-4$; (a): $\alpha=\pi/2$, $\varepsilon=0.1$; (c): $\alpha=\pi/2$, $\varepsilon=0.01$.
The singular basin of the adaptive phase oscillator (56) in Fig. 2(a,c) has a SF that extends into the region of negative $\mu$. This creates a channel through which initial conditions with arbitrary $\mu$ and selected $\varphi$ are attracted to the periodic rotation $\gamma_{c}$. Once again, this is impossible within the framework of the averaged system (3), where only initial conditions with $\mu>\mu_{b}$ can converge to the averaged periodic rotation $\gamma_{c}$; see Fig. 2(b).
Next, we provide details of how the boundaries of the SF in Fig. 2(a,c) are formed. As in the case of the adaptive pitchfork normal form (12), these boundaries consist of orbits attracted to the saddle equilibrium $e_{2}$ (branches of the stable manifold of $e_{2}$). Tracking these orbits backwards in time reveals that they pass close to the branch of unstable quasistatic equilibria of the fast subsystem (the dashed part of the curve $S$ in Fig. 2(a,c)), which is also known as the critical manifold [34]. Since this part of the critical manifold is attractive in backward time, the two SF boundaries exponentially approach each other as they are attracted to $S$, causing the SF to become exponentially narrow. After the SF extends past $S$ to negative $\mu$, it undergoes further rotations and repeatedly intersects the white basin of attraction of $e_{1}$ leading to a SF with a complicated pattern.
For the systems with a singular basin that have been considered so far, we can obtain the universal scaling of the SF volume in phase space as a function of $\varepsilon$ to estimate how quickly SFs shrink as $\varepsilon$ approaches 0. In all cases, the motion along the trajectories, which correspond to the SF boundaries, is slow and its speed is proportional to $\varepsilon$ along the unstable branch of a critical manifold $S$. Thus, the time they spend close to an unstable branch of $S$ of length $L$ scales as $L/\varepsilon$. The width $\delta$ of the SF is determined by two trajectories that pass close to $S$ (blue trajectories in Fig. 2). Hence, $\delta$ decreases exponentially as time goes backwards. Therefore, this width can be estimated as $\delta\sim\exp(-L\lambda/\varepsilon)$ at the moment when the SF leaves the neighbourhood of an unstable branch of $S$ of length $L$. Here $\lambda$ is an effective repulsion rate at $S$, which corresponds to the attraction in the reverse time. These arguments lead to the following expected scaling for the volume of a SF:
$V(\varepsilon)\sim\exp(-C\varepsilon^{-1}),\quad C>0.$ (8)
The scaling (8) is in perfect agreement with the adaptive normal form (12) and system (26). Furthermore, the adaptive phase rotator (56) exhibits this scaling too, as shown in Fig. 2(c). The SF scaling (8) is expected to hold for large classes of slow-fast systems, at least in low dimensions, as the arguments leading to (8) are based on the rather general geometric structure of the basin boundaries. However, as we show below, SFs with intricate geometries found in higher-dimensional systems may show deviations from this scaling.
Figure 3: (a) Bistable potential of the slow subsystem of two adaptive phase oscillators (9-10), for $N=2$, after adiabatic elimination and averaging of the fast variables $\varphi_{1,2}$. (b, d) Two cross-sections of the singular basin with a singular funnel (SF) of the same system: (b) $\varphi_{1}=1.2461$ and (d) $\mu=2.86$. The basin of attraction of the periodic rotation $\gamma_{c}$ is plotted in red and the basin of attraction of the stable equilibrium $e_{1}$ is plotted in white. The (blue) boundary of the two basins is given by the stable manifold of the saddle equilibrium $e_{2}$. (c) The volume of the SF limited to $10\leq\mu\leq 0$ as a function of $\varepsilon$ obtained using Monte Carlo simulation for different values of the parameter $\Delta_{\omega}=\omega_{1}-\omega_{2}$, showing the robustness of the SF. Other parameters: $\omega_{1}=-4$, $\kappa=1$, $\eta=10$; (b-d) $\varepsilon=0.1$.
A higher-dimensional class of systems, in which singular basins can be observed, is the mean-field coupled active rotators
$\dot{\varphi}_{i}$ $=\omega_{i}+\mu-\sin\varphi_{i}+\frac{\kappa}{N}\sum_{j=1}^{N}\sin(\phi_{j}-\phi_{i}),$ (9)
$\dot{\mu}$ $=\varepsilon(-\mu+\eta(1-X)),$ (10)
where $i=1,\dots,N$ and $X=\frac{1}{N}\sum_{j=1}^{N}\sin\left(\varphi_{j}+\alpha\right)$. In this system, the frequencies of the individual rotators are adapted globally by the slow variable $\mu$, and $\mu$ is driven by the mean-field $X$. In contrast to the extensively studied system of coupled rotators without adaptation [41, 36, 13, 29, 47, 7, 22, 48], the adaptive system (910) evolves on two distinct timescales with ratio $\varepsilon$. Adaptive systems of a similar nature have been shown to exhibit distinct dynamical properties such as canard cascading [10, 4], emergent excitability [8], and others [51, 9, 53]. Here, we add singular basins to this list of distinct properties.
Figure 3 shows singular basin and its scaling in system of $N=2$ rotators. The system features the coexistence of two attractors: one stable equilibrium and one periodic or quasi-periodic rotation. A bistable potential calculated by averaging system (9) with $N=2$ oscillators is shown in Figure 3(a). Two-dimensional cross-sections of the corresponding singular basin for fixed values of $\varphi_{1}$ and $\mu$ values are illustrated in Figs. 3(b) and (d), respectively. Figure 3(c) shows how the SF volume shrinks with $\varepsilon$ for different values of the detuning parameter $\Delta_{\omega}=\omega_{1}-\omega_{2}$. We observe that $V(\varepsilon)$ follows the scaling (8) for some values of $\Delta_{\omega}$ (e.g. $\Delta_{\omega}=0.6$), while for other values of $\Delta_{\omega}$, it exhibits resonance-like deviations from (8) (e.g. $\Delta_{\omega}=1$ or $\Delta_{\omega}=2.2$). Hence, we conclude that higher-dimensional systems can have SFs that deviate from the simple scaling (8) and leave this as a question for future research.
Figure 4: (a) The volume of the basin of attraction of ten adaptive phase oscillators (9-10) for $N=10$, limited to $10\leq\mu\leq 0$, as a function of $\varepsilon$ obtained using Monte Carlo simulation see the supplemental material fro more details. Its decrease with $\varepsilon$ indicates the presence of a singular funnel (SF). (b,c) Two trajectories starting from close initial conditions: $\varphi_{i}(0)=6$ for $i=1,\dots,10$, with $\mu(0)=-5$ (red) and $\mu(0)=-5.1$ (blue). The red trajectory starts from SF and therefore it is attracted to a stable rotating state. Other parameters: $\omega_{i}=-4+(i-1)/9$, $i=1,\dots,10$, $\eta=10$, $\alpha=\pi/2$, and $\kappa=1$.
A visualization of SFs in high-dimensional systems via two-dimensional cross sections is very difficult. Instead, the estimated volume of the SF for the case of $10$ coupled rotators (eleven-dimensional phase space) is plotted in Figure 4(a). Similar to the lower-dimensional examples, the volume shrinks with $\varepsilon$, indicating the presence of a SF. Additionally, Fig. 4(b) shows two trajectories: (red) one starting from the SF and converging to periodic rotations, and (blue) one starting and staying very close by, but eventually converging to the stable equilibrium $e_{1}$, which is another indication of a SF.
Finally, singular basins are a robust phenomenon. This is due to the fact that its ingredients persist under small parameter changes. This applies to the stable manifolds of the saddle equilibria and the coexisting attractors. We illustrate the parameter region of the singular basin and the possible mechanisms of its destruction for system (56) in the supplemental material.
In summary, we have uncovered singular basins with singular funnels (SFs) in bistable slow-fast systems, and showed that they are robust and ubiquitous. The SFs prevent useful and widely used dimensionality reductions, even when the timescales differ significantly and there is a unique stable quasistatic state that changes continuously with the slow variable(s). Although adiabatic elimination or averaging of the fast variables retains the coexistence of different stable states and different basins of attraction in the reduced slow subsystem, the fast dynamics of the full system can lead to transitions between different stable states via SFs that are not possible in the slow subsystem. From the slow subsystem’s perspective, these unexpected dynamics resemble quantum-like ‘tunnelling’ or ‘teleporting’ from one attractor to another. We have constructed low-dimensional canonical systems that can be considered normal forms for a singular basin and showed a universal scaling of SFs with increasing timescale ratio in these systems. Additionally, we have demonstrated that SFs can form intricate structures in the phase space of higher-dimensional systems with slowly adaptive mean-field coupling. A challenge for future research is to understand singular basins in systems with more than two timescales and more than two coexisting stable states.
Perhaps most intriguingly of all, any mathematical model of a real-world system is necessarily an approximation that omits some fast variables. Thus, our results also provide new insight into the potential for unexpected critical transitions [3]. For example, transitions between different stable states, which are not captured by the models due to omitting some fast variables, could be triggered in real systems.
Acknowledgements.
This work emanated from the research funded by Taighde Éireann—Research Ireland (Grant No. FFP-A/12066).
Supplemental Material
I Details of the numerical methods
I.1 The singular funnel volume using Monte Carlo simulations in Figures 2 and 3
We estimated the volume of the singular funnel in Figures 2 and 3 of the main paper by Monte Carlo simulations as follows:
  • For each value of $\varepsilon$, we randomly choose a large number $M$ of uniformly distributed initial conditions, with $\varphi_{i}(0)\in[0,\,2\pi)$ and $\mu(0)\in[-10,\,0]$.
  • For each of these initial conditions, we numerically solve the initial value problem consisting of system (5)–(6) or (9)–(10) and the initial condition, for a long enough time. For the numerical integration we use the MATLAB function ode45, with $\texttt{RelTol}=10^{-10}$ and $\texttt{AbsTol}=10^{-10}$. The time span for our integration is $0,\,10/\varepsilon]$.
  • We then determine whether the solution trajectory converges to the stable equilibrium $e_{1}$ or to the rotating periodic orbit $\gamma_{c}$. To detect convergence, we use a user-defined event function to stop the integration at time $t_{\mathrm{end}}$ when $\mu(t_{\mathrm{end}})>9$. If this condition applies, then the solution has converged to $\gamma_{c}$ otherwise, it has converged to $e_{1}$.
  • The volume $V$ of the basin of attraction of the limit cycle $\gamma_{c}$ is then given by
    $V=\frac{K}{M},$
    where $K$ is the number of initial conditions whose solution trajectories converge to $\gamma_{c}$ over time.
I.2 Numerical averaging in Figure 3
We computed the potential of the average system of the two oscillator network in Figure 3 as follows:
  • For fixed parameter values of $\omega_{1,2},\,\kappa,\,\eta,\,\alpha$, and a given value of $\mu$, we randomly choose initial values for $\varphi_{1}(0)$ and $\varphi_{2}(0)$ from the interval $0,\,2\pi)$.
  • We consider the (fast) layer problem of system (5)–(6), which is given as:
    $\begin{split}\dot{\varphi}_{i}=\omega_{i}+\mu-\sin\varphi_{i}+\frac{\kappa}{N}\sum_{j=1}^{N}\sin(\varphi_{j}-\varphi_{i}),\quad i=1,2.\end{split}$ (11)
  • The initial value problem consisting of system (11) and the randomly chosen initial condition $\left(\varphi_{1}(0),\,\varphi_{2}(0)\right)$ was solved numerically for a sufficiently long time interval [0, 600][0,\,600]. For the numerical integration, the MATLAB function ode45 was used, with $\texttt{RelTol}=10^{-10}$ and $\texttt{AbsTol}=10^{-10}$.
  • The transient part $t\in[0,100]$ of the solution was discarded to ensure that the solution has converged to an invariant state.
  • We compute $X(t)=\frac{1}{2}\sum_{j=1}^{2}\sin(\varphi_{j}(t)+\alpha)$ for $t\in[100,600]$ and write the average system as:
    $\dot{\mu}=\bar{g}(\mu)=\frac{1}{500}\int_{100}^{600}\left(-\mu+\eta(1-X(t))\right)\,dt$ (12)
    The definite integral on the right-hand side was computed numerically using the trapezoidal rule.
  • The potential $U(\mu)$ is given by:
    $U(\mu)=\int_{0}^{\mu}\bar{g}(s)\,ds.$
II Supercritical pitchfork normal form with adaptive parameter
The pitchfork normal form with adaptively changing parameter, as introduced in the manuscript, is given by
$\dot{x}(t)$ $=x(\mu-x^{2}),\quad x\geq 0,$ (13)
$\dot{\mu}(t)$ $=\varepsilon(-\mu+ax-b),$ (14)
where $\varepsilon>0$ is a small parameter, and $a,b>0$ are parameters determining the linear adaptation function.
First we consider the equilibria of the fast system (13), which, when considered as a set in the phase space $(x,\mu)$ of the fast-slow system (13)–(14), define the critical manifolds
$S_{0}$ $=\{(x,\mu)\in[0,\infty)\times\mathbb{R}\ :\ x=0\},$ (15)
$S_{1}$ $=\{(x,\mu)\in[0,\infty)\times\mathbb{R}\ :\ x=\sqrt{\mu},\ \mu\geq 0\}.$ (16)
Linearization of the vector field along these manifolds shows that $S_{0}$ has two branches, namely
$S_{0}^{\textup{a}}$ $=\{(x,\mu)\in[0,\infty)\times\mathbb{R}\ :\ x=0,\,\mu<0\},$ (17)
$S_{0}^{\textup{r}}$ $=\{(x,\mu)\in[0,\infty)\times\mathbb{R}\ :\ x=0,\,\mu>0\},$ (18)
which are locally attracting and repelling respectively and such that $S_{0}=S_{0}^{\textup{a}}\cup\left\{0\right\}\cup S_{0}^{\textup{r}}$, while $S_{1}$ is locally attracting. This stability information is shown in Fig. 5(b) by solid lines for the stable parts and dashed line for the unstable part of the critical manifolds respectively.
Thus the fast system (13) has a unique stable equilibrium for all values of the slow variable $\mu$, which is attracting all values of $x\geq 0$ for $\mu\leq 0$ and all $x>0$ for $\mu>0$. Therefore we consider the reduced slow system on the union of the corresponding stable parts of the critical manifolds. We substitute
$x=\sqrt{\mu H(\mu)}=\begin{cases}0,&\mu\leq 0\\ \sqrt{\mu},&\mu>0,\end{cases}$ (19)
into (14) and obtain
$\dot{\mu}=\varepsilon\begin{cases}-\mu-b,&\mu\leq 0\\ -\mu+a\sqrt{\mu}-b,&\mu>0,\end{cases}$ (20)
or in a more compact form
$\dot{\mu}=\varepsilon\left(-\mu+a\sqrt{\mu H(\mu)}-b\right),$ (21)
where $H(\mu)$ is the Heaviside step function.
Figure 5: Phase portrait of the pitchfork normal form, system (13)–(14). (a) $\varepsilon=0.1$, the red region (white) region is the basin of attraction of $e_{0}(e_{2})$, the stable $W^{s}(e_{1})$ (unstable $W^{u}(e_{1})$) manifolds of the saddle point $e_{1}$ in blue. (b) $\varepsilon\to 0$, critical manifolds $S_{0,1}$ in thick black, the attracting parts in solid and the repelling part of $S_{0}$ is dashed. The double arrowed vertical black lines indicate the direction of the fast layer system. Parameters: a=3,b=2a=3,\,b=2.
We consider the case when the reduced system (21) has three equilibria $e_{0}$, $e_{1}$, and $e_{2}$ as shown in Fig. 5, their $\mu$ coordinates are $\mu_{0}<0$, $\mu_{1}>0$, and $\mu_{2}>0$, respectively. Requiring three equilibria leads to the following conditions on the parameters:
$b>0\quad\text{and}\quad a>2\sqrt{b}.$ (22)
Under the conditions (22), the reduced system (21) is bistable with two stable equilibria $e_{0}$, $e_{2}$ and one unstable equilibrium $e_{1}$:
$e_{0}:(0,-b),\quad e_{1,2}:\left(\sqrt{\mu},\frac{a}{2}\pm\sqrt{\left(\frac{a}{2}\right)^{2}-b}\right).$ (23)
The potential of system (21) can be calculated as the integral of its right-hand side, leading to
$U(\mu)=\frac{\mu^{2}}{2}+b\mu-\frac{2}{3}a\mu\sqrt{\mu H(\mu)},$ (24)
where, for simplicity, we dropped the scaling factor $\varepsilon$.
The equilibria $e_{0}$ and $e_{2}$ are stable in the phase space of the full system (13)–(14). The attraction basins for these equilibria are separated by the stable manifolds $W_{1,2}^{s}(e_{1})$ of the saddle equilibrium $e_{1}$, see Fig. 5 and Fig. 1(a) of the main manuscript. For our purposes, it is worth noting that the manifold $W_{2}^{s}(e_{2})$ converges exponentially to $x=0$ as $\mu\to\infty$. As a result, a part of the attraction basin of the equilibrium $e_{0}$ extends into an exponentially small region for all $\mu>0$, which we call singular funnel.
III A model without critical manifold crossing
Another system of equations introduced in the manuscript is
$\dot{x}(t)$ $=x\left(\tanh\mu+2-x\right),$ (25)
$\dot{\mu}(t)$ $=\varepsilon(-\mu+ax-b).$ (26)
The corresponding reduced system on the stable critical manifold (adiabatic elimination) is given by
$\dot{\mu}(t)$ $=\varepsilon(-\mu-b+a\left(\tanh(\mu)+2)\right),$ (27)
and the corresponding (rescaled) potential $V$ is
$U$ $=-\int\left(a\tanh(\mu)-\mu-b+2a\right)\,d\mu$
$=\mu^{2}/2-(2a-b)\mu-a\ln(\cosh(\mu)).$
IV Adaptive phase rotator
IV.1 Timescale reduction
The adaptive phase rotator considered in the manuscript has the form
$\dot{\varphi}$ $=\omega+\mu-\sin\varphi,$ (28)
$\dot{\mu}$ $=\epsilon(-\mu+\eta(1-\sin(\varphi+\alpha))).$ (29)
The slow-fast dynamics of this system for $\alpha=0$ was studied in [18], for the deterministic and stochastic case. Here we extend these results (for the deterministic case) to $\alpha\neq 0$.
The critical manifold has the form:
$S_{0}:=\left\{(\mu,\varphi):\ \mu=-\omega+\sin\varphi\right\},$ (30)
and it exists in the stripe $|\mu+\omega|\leq 1$. To obtain the dynamics on this critical manifold (adiabatic elimination), we substitute $\sin\varphi=\mu+\omega$ into (29), leading to
$\dot{\mu}=-\mu+\eta\left(1-\left(\mu+\omega\right)\cos\alpha\mp\sqrt{1-\left(\mu+\omega\right)^{2}}\sin\alpha\right),\quad\left|\mu+\omega\right|\leq 1.$ (31)
For $\left|\mu+\omega\right|>1$, the fast subsystem (28) has no equilibria, but exhibits periodic rotation described by
$\varphi_{\mu}(t)=2\arctan\left(\frac{1+\Omega(\mu)\tan\left(\frac{t}{2}\Omega(\mu)\right)}{\omega+\mu}\right),$
where
$\Omega(\mu)=\sqrt{\left(\omega+\mu\right)^{2}-1}.$
To average the slow dynamics (29) along these fast rotations, we need to average the oscillating term $\sin(\varphi(t)+\alpha)$ over the period $T=2\pi/\Omega$:
$\left\langle\sin(\varphi_{\mu}(t)+\alpha)\right\rangle=\frac{1}{T}\int_{0}^{2\pi/\Omega}\sin\left(\varphi_{\mu}(t)+\alpha\right)dt=$
$\frac{1}{T}\cos\alpha\int_{0}^{T}\sin\varphi_{\mu}(t)dt+\frac{1}{T}\sin\alpha\int_{0}^{T}\cos\varphi_{\mu}(t)dt=$
$\left(\left(\omega+\mu\right)-\Omega\right)\cos\alpha+\frac{1}{T}\sin\alpha\int_{0}^{T}\cos\varphi_{\mu}(t)dt$
$\left(\left(\omega+\mu\right)-\Omega\right)\cos\alpha+\frac{1}{T}\sin\alpha\int_{0}^{2\pi}\frac{\cos\varphi d\varphi}{\omega+\mu-\sin\varphi}=$
$\left(\left(\omega+\mu\right)-\Omega\right)\cos\alpha+\frac{1}{T}\sin\alpha\int_{0}^{2\pi}\frac{d\sin\varphi}{\omega+\mu-\sin\varphi}=$
$\left(\left(\omega+\mu\right)-\Omega\right)\cos\alpha.$
Therefore, the averaged equation is
$\dot{\mu}=-\mu+\eta\left(1-\left(\left(\omega+\mu\right)-\sqrt{\left(\omega+\mu\right)^{2}-1}\right)\cos\alpha\right),\quad|\omega+\mu|>1.$ (32)
Finally, we combine the reduced systems (31) for $|\mu+\omega|\leq 1$ and (32) for $|\mu+\omega|>1$, to obtain
$\dot{\mu}=-\mu+\eta\left(1-\left(\mu+\omega\right)\cos\alpha+\Omega(\mu)\right),$ (33)
where
$\Omega(\mu)=\begin{cases}\sqrt{1-\left(\mu+\omega\right)^{2}}\sin\alpha,\quad\left|\mu+\omega\right|\leq 1\\ \sqrt{\left(\omega+\mu\right)^{2}-1}\cos\alpha,\quad\left|\mu+\omega\right|>1.\end{cases}$ (34)
The equilibria of the averaged dynamics satisfy
$\mu+\eta\left(1-\left(\mu+\omega\right)\cos\alpha+\Omega(\mu)\right)=0.$ (35)
IV.2 Parameter region for singular basin
In Figure 6, we present the bifurcations of system (28)–(29) with respect to the parameters $\alpha$ and $\varepsilon$. In the parameter region we examined, there are two Hopf bifurcation curves and three homoclinic curves. Also, the system has two equilibrium solutions $e_{1}$ and $e_{2}$, and up to four periodic solutions: $\gamma_{1,2}$, which are regular limit cycles, and $\gamma_{c,u}$, which are rotating limit cycles resulting from the fact that $\varphi\in[0,2\pi)$. We point out that the rotating limit cycle $\gamma_{c}$ is always stable in the parameter region under examination.
If we consider Figure 6 (a), starting from the left-hand side, we have a monostable system in region (b), Figure 6 (b), where the equilibrium $e_{1}$ is unstable. At the subcritical Hopf curve $H_{1}$, the unstable equilibrium $e_{1}$ gains stability in region (c), Figure 6 (c), and an unstable limit cycle $\gamma_{1}$ emerges to form the boundary of the basin of attraction of $e_{1}$.
The unstable limit cycle $\gamma_{1}$ intersects the saddle equilibrium $e_{2}$ at the homoclinic bifurcation $h_{1}$, Figure 6 (d). To the right-hand side of $h_{1}$, region (g), the basin of attraction of the rotating limit cycle is singular, with the boundary given by the stable manifold of the saddle $e_{2}$, Figure 6 (g).
The second homoclinic bifurcation, $h_{2}$, is formed by the intersection of the unstable rotating limit cycle $\gamma_{u}$ and the saddle equilibrium $e_{2}$, Figure 6 (g). Region (e), Figure 6 (e), below the curve $h_{2}$, is a bistable region where both the equilibrium $e_{1}$ and the rotating limit cycle $\gamma_{c}$ are stable, and the basin boundary is given by the unstable rotating limit cycle $\gamma_{u}$. The second Hopf bifurcation, $H_{2}$, is supercritical. The stable equilibrium $e_{1}$ in Figure 6 (g) loses stability, giving rise to a stable limit cycle $\gamma_{2}$, Figure 6 (h). The basin of attraction of $\gamma_{c}$ is still singular in region (h). The region of singular basin ends at the homoclinic bifurcation $h_{3}$, Figure 6 (i), where the stable limit cycle $\gamma_{2}$ intersects the saddle equilibrium $e_{2}$. In region (j), Figure 6 (j), the system is monostable again, where the rotating limit cycle $\gamma_{c}$ is the only stable attractor.
Figure 6: Two-parameter $(\alpha,\,\varepsilon)$ bifurcation diagram for system (28)–(29), with examples of phase portraits.
IV.3 Analytical considerations for the singular funnel scaling
Here we provide some analytical considerations, which substantiate the scaling given in the main manuscript as equation (8).
Consider $\varepsilon\ll 1$ and the boundaries of the singular basin to be
$\phi_{u}(t)=\phi(t,\phi_{u0},\mu_{0}),\quad\mu_{u}(t)=\phi(t,\phi_{u0},\mu_{0})$
for the upper, and
$\phi_{l}(t)=\phi(t,\phi_{l0},\mu_{0}),\quad\mu_{l}(t)=\phi(t,\phi_{l0},\mu_{0})$
for the lower boundaries. Here $\phi_{l0}<\phi_{u0}$, $\omega-1<\mu_{0}<\mu(e_{2})$ are some chosen points on the singular funnel such that $\phi(0,\phi_{X0},\mu_{0})=\phi_{X0}$ and $\mu(0,\phi_{X0},\mu_{0})=\mu_{0}$ for $X\in\{u,l\}$. For any fixed $\delta<\phi_{u0}-\phi_{l0}$, there exists $t_{0}<0$ such that
$\phi_{l}(t_{0})-\phi_{u}(t_{0})=\delta$
due to the contraction to the equilibrium of the layer system (fast equation (28)) in the backwards time.
Being $\delta$-close to the equlibrium of the fast system, it can be linearized for $t_{f}<t<t_{0}$, where $t_{f}$ is the time when the system approaches the fold point of the layer equation. This leads to
$\dot{\phi}$ $=\omega+\mu-\sin(\phi^{*}(\mu))-\cos(\phi^{*}(\mu))\left(\phi-\phi^{*}(\mu)\right),$
$\dot{\mu}$ $=\epsilon\left(-\mu+\eta\left(1-\sin(\phi^{*}(\mu)+\alpha)-\cos(\phi^{*}(\mu)+\alpha)\right)\left(\phi-\phi^{*}(\mu)\right)\right).$
Taking into account that $\phi^{*}(\mu)$ is the critical manifold, the leading terms in both equations read as:
$\dot{\phi}$ $=-\cos(\phi^{*}(\mu))\left(\phi-\phi^{*}(\mu)\right),$
$\dot{\mu}$ $=\epsilon\left(-\mu+\eta\left(1-\sin(\phi^{*}(\mu)+\alpha)\right)\right).$
Denoting $\Delta=\phi_{u}-\phi_{l}$, we have
$\dot{\Delta}(t)$ $=-\cos(\phi^{*}(\mu))\Delta(t),$
$\dot{\mu}(t)$ $=\epsilon\left(-\mu+\eta\left(1-\sin(\phi^{*}(\mu)+\alpha)\right)\right).$
The above equations give $d\Delta/dt$ and $d\mu/dt$. Hence we obtain $d\Delta/d\mu$ as follows
$\frac{d\Delta}{d\mu}=\frac{1}{\epsilon}\frac{-\cos(\phi^{*}(\mu))}{-\mu+\eta\left(1-\sin(\phi^{*}(\mu)+\alpha)\right)}\Delta,$
which can be solved as
$\Delta(\mu)=\Delta(\mu_{0})\exp\left[\frac{1}{\epsilon}\int_{\mu_{0}}^{\mu}\frac{\cos(\phi^{*}(\mu))}{\mu-\eta\left(1-\sin(\phi^{*}(\mu)+\alpha)\right)}d\mu\right]=\Delta(\mu_{0})\exp\left[-\varepsilon^{-1}C(\mu,\mu_{0})\right],$ (36)
which provides the scaling as in equation (8) of the main paper.
In addition, according to [31], there is only an algebraic contraction of the singular funnel stripe across the fold point. Hence, the exponential estimate (36) remains for the further motion along the fast flow.